Subshilava:

Someone please explain this I keep getting -2x^2-12x-16 But that’s not right 😭 (Pic below)

Subshilava:

1 attachment
Ash2746:

What grade are you in? in you don't mind me asking

Subshilava:

11th I have the function-2(x+4)(x-2) but I cannot get how to write it in standard form 😪

Subshilava:

Like how do they have -2(x^3

Ash2746:

I thought so that was about to hurt my brain for a moment

Subshilava:

Idk they have a video but the lady never put the degree until she gets to graphing in just confused on where they get the x^3

sllo:

polynomial is the highest power of x, which is 3

Subshilava:

Actually I sort of figured out what I was doing wrong I was multiplying instead of writing it multiple times 🤦♀️

Subshilava:

Nvm I still can’t get it

Subshilava:

@hero wrote:
@Subshilava, what is the result of factoring -2 out of the expression?
Idk I get to the end of multiplying and the numbers just mix together she used the box method is there n easier method where the numbers don’t mix together? But for the factor form question I did end up with -2(x+2)^2(x-1)^2

sllo:

@subshilava wrote:
@hero wrote:
@Subshilava, what is the result of factoring -2 out of the expression?
Idk I get to the end of multiplying and the numbers just mix together she used the box method is there n easier method where the numbers don’t mix together? But for the factor form question I did end up with -2(x+2)^2(x-1)^2
You've already factored the function as -2(x+2)^2(x-1)^2. This means you have two binomials raised to the power of 2. Expand (x+2)^2 and (x-1)^2, (x+2)^2 = x^2 + 4x + 4 (x-1)^2 = x^2 - 2x + 1 then sub and multiply, -2(x^2 + 4x + 4)(x^2 - 2x + 1) Distribute and simplify multiplying and combining like terms, you should get: -2x^4 + 4x^3 + 12x^2 - 16x - 8 I think thats the answer

Hero:

@sllo wrote:
@subshilava wrote:
@hero wrote:
@Subshilava, what is the result of factoring -2 out of the expression?
Idk I get to the end of multiplying and the numbers just mix together she used the box method is there n easier method where the numbers don’t mix together? But for the factor form question I did end up with -2(x+2)^2(x-1)^2
You've already factored the function as -2(x+2)^2(x-1)^2. This means you have two binomials raised to the power of 2. Expand (x+2)^2 and (x-1)^2, (x+2)^2 = x^2 + 4x + 4 (x-1)^2 = x^2 - 2x + 1 then sub and multiply, -2(x^2 + 4x + 4)(x^2 - 2x + 1) Distribute and simplify multiplying and combining like terms, you should get: -2x^4 + 4x^3 + 12x^2 - 16x - 8 I think thats the answer
Incorrect.

sllo:

@hero wrote:
@sllo wrote:
@subshilava wrote:
@hero wrote:
@Subshilava, what is the result of factoring -2 out of the expression?
Idk I get to the end of multiplying and the numbers just mix together she used the box method is there n easier method where the numbers don’t mix together? But for the factor form question I did end up with -2(x+2)^2(x-1)^2
You've already factored the function as -2(x+2)^2(x-1)^2. This means you have two binomials raised to the power of 2. Expand (x+2)^2 and (x-1)^2, (x+2)^2 = x^2 + 4x + 4 (x-1)^2 = x^2 - 2x + 1 then sub and multiply, -2(x^2 + 4x + 4)(x^2 - 2x + 1) Distribute and simplify multiplying and combining like terms, you should get: -2x^4 + 4x^3 + 12x^2 - 16x - 8 I think thats the answer
Incorrect.
Very correct

sllo:

Wait no its -2(x+2)^2(x-1)^2.

Subshilava:

@hero wrote:
Also @subshilava, you shouldn't post a problem without also posting the original and complete instructions. The way you posted this problem is improper. You're preventing people from being able to help you properly.
That’s all that was provided is what was in the image I just needed help with the writing standard form part

Subshilava:

Also the only 3 is the degree

Hero:

So, I figured out the correct expression in standard form. First off, the original expression according to the properties given is \(-2(x-1)^2(x+2)\). This, when expanded becomes \(-2(x^3-3x+2)\) which is in the standard form you were looking for.

Hero:

@Subshilava

Hero:

@sllo you almost had it correct, but only one of the factors has multiplicity 2.

Subshilava:

Oh ok i wonder why it must’ve been a mistake on the teachers side

Hero:

@subshilava wrote:
Oh ok i wonder why it must’ve been a mistake on the teachers side
I assumed that at first. There's no mistake on the instructor's side. The correct expression is \(-2(x+2)(x-1)^2\) and you now have the correct standard form.

Hero:

But the instructor should not have put the characteristics of the function at the very bottom.